MIT/CXC Group Meeting, 2 Aug 1999
THE
NOTES
Spacecraft news (CRC)
- Chandra operations going well.
Duty Scientist news (DPH, et al.)
- Automatic processing working. Some
warnings and errors, but nothing serious. Some CC3x3 problems being
worked.
Last mtg actions:
- Region, psf progress checks (DPH): still pending.
- Poster printer (DPH): Bautz agreed to share cost. DPH to get quotes.
- Grating RMF, ARF (DPH, DSD, CB, MW) : in progress.
- Barycentric correction (DPH): not done in Level 1. Left to user on
per-source basis, since some applications the small time
difference over the field may matter. Software will come from
Arnold Rots.
HETG insertion review (DD)
- Talked to people ---too busy to meet. To
check: what happens if insertion command given when grating is already
in? (issue is coordination/knowledge of off-line and on-line commands).
ACIS news (various)
- CC-3x3 ran through Level 1. Checks revealed that x,y were
flipped.
- Local area set up for Level 1 products. Beware of bogus keywords! (in
manual processing, correct observational parameters are not propagated
into the header.
- Gain change vs CCD node from cal source: S3 is different from ground cal.
- There is a purported SIM cal source for ACIS when HRC-S at aim-point.
Where'd that come from? What's it's purpose?
S/C issues (CRC)
- The Slosh: There is extra IPS fuel since the thrusters are very well
aligned, and fewer residual corrections were needed. Slosh could
affect resolution (spatial and spectral). Empirical tests indicate
very small affect. More tests are planned.
TG Analysis issues (DD)
- What are the default binning criteria? Current defaults give 6972
bins. Why? Because the default ranges were chosen to allow some
variance in aim-point, and cover, e.g., 1-40A of MEG bins at about 2x
oversampling. This needs to be reviewed for optimum "nice" defaults
(e.g., a power of 2, and LETG/HRC-S a multiple of that?). This is
important for producing a standard set of ARFs and RMFs. DPH to review.
XRCF Data Analysis (HLM)
- Chip positions: get 3-4 pixel errors in tiled detector coordinate gaps
(TDET). Determined x-corrections from data. Need to propagate back to
the data system (pixlib, cal data, grating coordinates). Need to
review ground measurements of as-built ACIS. (Action to DPH, DD.)
- Efficiency modeling: broad bump near W-Mbeta line; 2.2 keV edge
feature? Due to gold M-edge model? Size is 5-7%.
- Ratios of MEG:HEG:LEG efficiencies give 5-10% residuals in
well-exposed regions.
Emission Line Project (CRC)
- Plan meeting here with interested MSFC people.
Federal funding (CRC)
- The House Appropriations Cte. has slashed the NASA
budget. Budget goes to the floor this week for full House vote.
Senate proposes their budget later. Letters to your representatives
are needed!
"It's Working" HETG papers (HLM)
- Papers should be drafted. E.g., Capella data
will come soon. Draft outlines of papers.
ACTIONS:
(explicit or implied; NEW actions get a "*")
- * DPH: Draft outline of Capella first-hetg-spectrum paper (due
Friday 8/6).
- * JH: organize Emission Line project meeting.
- * DPH: review default spectral binning parameters; revise if
necessary.
- * DPH,DD: Investigate pixlib/HAK changes required for chip-gap
discrepancies.
- DPH: investigate region support/development status in CXC
visualization and tools.
- ALL: Review GTO schedule (NSS to send URL).
- DPH: do Capella simulation, process to L2 (HETG and LETG). (in
progress: source models made, simulations in progress)
- DPH: pursue poster printer purchase. (open)
- DPH, MW, DSD: get ARFs and RMFs on-the-table. (open)
Next meeting: Wed 11 Aug 1999, 2:00p
dph@space.mit.edu